OpenStudy (randy):

x-0.09x=728 need to see how to solve

OpenStudy (anonymous):

0.91x =728 so x = 728/0.91 = 800 as (91x8=728)

OpenStudy (anonymous):

\[x-0.09x = 728\] you have on the left side two like variables so combine them the first \[x = 1*x\] so \[1-0.09 = 0.91\] you then have \[0.91x = 728\] then x = 728/0.91